40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 8 06:29:06 2011, kiwamu, Update, Green Locking, Intensity noise setup 
    Reply  Wed Mar 9 04:46:13 2011, kiwamu, Update, Green Locking, more intensity noise measurement intensity_setup.pngIntensityNoise.png
       Reply  Wed Mar 9 18:17:11 2011, kiwamu, Update, Green Locking, Intensity noise coupling diff_noise.pngIN_TF.png
          Reply  Thu Mar 10 14:06:54 2011, kiwamu, Update, Green Locking, Intensity noise limits the beatnote sensitivity diff_noise_Mar8.pngIN_TF_complete.png
             Reply  Thu Mar 10 14:22:58 2011, kiwamu, Update, Green Locking, Intensity noise limits the beatnote sensitivity 
                Reply  Thu Mar 10 14:29:05 2011, Koji, Update, Green Locking, Intensity noise limits the beatnote sensitivity 
                Reply  Thu Mar 10 14:30:53 2011, rana, Update, Green Locking, Intensity noise limits the beatnote sensitivity 
                   Reply  Thu Mar 10 16:00:53 2011, Aidan, Joe, Update, Green Locking, Intensity stabilization loop using beatnote DC. P1000313.jpgP1000314.jpgP1000315.jpgGREEN_ISS_LOOP.pdfScreenshot-C1ALS_OVERVIEW.adl.png
                      Reply  Thu Mar 10 21:45:34 2011, rana, Update, Green Locking, Intensity stabilization loop using beatnote DC. 
Message ID: 4383     Entry time: Tue Mar 8 06:29:06 2011     Reply to this: 4389
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Intensity noise setup 

[Jenne, Chris, Kiwamu]

 A photo diode and an AOM driver have been newly setup on the PSL table to measure the intensity noise coupling to the beat note signal.

We tried taking a transfer function from the PD to the beat, but the SNR wasn't sufficient on the PD. So we didn't get reasonable data.

 

(what we did)

  - put a DCPD after the doubling crystal on the PSL table. The PD is sitting after the Y1 mirror, which has been used for picking off the undesired IR beam.

  - installed the AOM driver (the AOM itself had been already in place)

  - injected some signals onto the AOM to see if we can see an intensity fluctuation on the PD as well as the beat signal

 

(intensity noise)

  In order to have better SNR for the intensity measurement, we put an AC coupled SR560 with the gain of 100 just before the ADCs.

When a single frequency signal was applied from a Stanford Research's function generator to the AOM, we could clearly see a peak at the doubled frequency of the injected signal.

Also a peak at the same frequency was found on the beat note signal as well.

But when random noise was injected from the same function generator, the random noise looked below the ADC noise.

Jenne adjusted the output voltage from the PD to about 1 V to avoid a saturation in the analog path, but later we realized that the ADC counts was marely ~ 20 counts.

So we will check the ADC tomorrow. Hopefully we will get a good SNR.

ELOG V3.1.3-